1) There are 7 flights daily from Newark International Airport to Bermuda.  Suppose the probability that any...

1)

There are 7 flights daily from Newark International Airport to Bermuda.  Suppose the probability that any flight arrives late is 0.10.  

What is the probability that more than 2 flights are late today?

a.

0.026

b.

.478

c.

.150

d.

.372

2)

There are 7 flights daily from Newark International Airport to Bermuda.  Suppose the probability that any flight arrives late is 0.10.  

What is the probability that five of the flights are late today?  (use the excel sheet to help answer the question)

a.

0.000

b.

0.023

c.

.372

d.

0.124

3)

There are 7 flights daily from Newark International Airport to Bermuda.  Suppose the probability that any flight arrives late is 0.10.  

What is the probability that none of the flights are late today?  (use the excel sheet to help answer the question)

a.

0.124

b.

.478

c.

0.023

d.

.372
Answer #1

1)

here this is binomial with parameter n=7 and p=0.1
P(X>2)=1-P(X<=2)= 1-∑x=0x-1   (nCx)px(q)(n-x) = 0.026

2)

P(X=5)= (nCx)px(1−p)(n-x)    = 0.000

3)

P(X=0)= (nCx)px(1−p)(n-x)    = 0.478